A helium balloon has a volume of 0.75 L. A student wants to release the balloon and have it go through a hole in a suspended plastic sheet. In order to do this, she must reduce the volume of the balloon to 0.69 L. If the temperature of the room is 22 degrees Celsius, what must she lower the temperature to so the balloon will fit through the hole? *

a) 16 degrees Celsius

b) 20 degrees Celsius

c) -14 degrees Celsius

d)-1.6 degrees Celsius
1
Expert's answer
2020-06-25T07:52:15-0400

V1/T1=V2/T2;

T2=V2*T1/V1=0.69*295/0.75=271.4 K=-1.6 degrees Celsius

Answer: d)